OpenStudy (anonymous):

partial fractions to evaluate: (x^4+3x^3+2x^2+1)/ (x^2+3x+2)

myininaya (myininaya):

first do long division since degree of numerator> degree of denominator

myininaya (myininaya):

x^2 ___________________________ x^2+3x+2 | x^4 +3x^3 +2x^2 +1 -(x^4 +3x^3 +2x^2) -------------------------- 1 \[\frac{x^4+3x^3+2x^2+1}{(x+2)(x+1)}=x^2+\frac{1}{(x+2)(x+1)}\]

myininaya (myininaya):

now you use partial fractions for that fraction write that \[\frac{1}{(x+2)(x+1)}=\frac{A}{x+2}+\frac{B}{x+1}\]

myininaya (myininaya):

\[1=A(x+1)+B(x+2)\] \[1=x(A+B)+(A+2B)\]

myininaya (myininaya):

=> A+B=0 =>A+2B=1 first equation=> A=-B pluggin' this into second equation we have -B+2B=1 B=1 but remember A=-B so A=-(1)=-1 so we have that \[\frac{x^4+3x^3+2x^2+1}{x^2+3x+2}=x^2+\frac{-1}{x+2}+\frac{1}{x+1}\]

OpenStudy (anonymous):

THANK YOU !